OpenStudy (anonymous):

Mark any solutions to the equation tan2x - 3 = 0. A. 2π/3 B. π C. -π/3 D. 5π/3

OpenStudy (anonymous):

oops that's suppose to be tan^2 x - 3 = 0

OpenStudy (anonymous):

@blanderos24

OpenStudy (anonymous):

Hum, this equation is \[\tan(2x-3)=0\] right?

OpenStudy (anonymous):

No its \[\tan ^{2}-3=0\]

OpenStudy (anonymous):

\[\tan^2x-3=0\]

OpenStudy (anonymous):

OK. \[\tan^{2}x-3=0\] \[(\tan x+\sqrt{3})(\tan x-\sqrt{3})=0\] \[\tan x=\pm \sqrt{3}\] \[x=\pm \frac{ \pi }{ 3 }\] You select the positive value. So answer is C.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

oops bumbed the wrong question